OpenStudy (anonymous):

Use a ratio to compare the values. Simplify the ratio. 5 m to 100 cm. HELP PLEASE!!

OpenStudy (imstuck):

What exactly do you need to do with this? Convert to m or cm? You need to be specific with the instructions.

OpenStudy (anonymous):

thats the problem 5 m to 100 cm make it a ratio and simplify

OpenStudy (anonymous):

5m = 500 cm

OpenStudy (anonymous):

So I would say 5:1 is answer

OpenStudy (anonymous):

Because it really 500/100

OpenStudy (anonymous):

thanks Shay!

OpenStudy (anonymous):

I can always count of u !

OpenStudy (anonymous):

on lol

OpenStudy (imstuck):

Ok, then you need to convert to either cm or m. We will do both! Set it up like this:\[5m \times \frac{ 100cm }{ 1m }\] to solve for cm. The meters cancel out and you multiply 5 by 100 to get 500cm. 500 cm/100 cm = 5/1 or 5:1. This is going from larger to smaller. If you went smaller to larger, it would be 100 cm/500 cm = 1/5 or 1:5. The other one would convert like this:\[100cm \times \frac{ 1m }{ 100cm }\] When the cm cancel out you are left with 100/100 which = 1. So the ratio here is 1m/5m or 1:5. Or the other way it is 5m/1m or 1:5. Hope the explanation helps, not just giving you the answer with no reason.

OpenStudy (anonymous):

Yw !

OpenStudy (anonymous):

wow thanks for that long explanation!

OpenStudy (anonymous):

It is definitely 5:1

OpenStudy (anonymous):

@Lindsay_552 don't choose 1:5 it's wrong

OpenStudy (anonymous):

5 m to 100 cm in simplified ratio is 5:1. This is calculated by first converting 5 m to cm so as to have common units. Therefore, 1 m = 100 cm, 5 m = 5*100 = 500 cm. 500 is to 100 is then simplified to 5:1.

OpenStudy (anonymous):

@IMStuck you are confused

OpenStudy (anonymous):

I picked 5.1

OpenStudy (anonymous):

ok good @Lindsay_552

OpenStudy (anonymous):

lol